OpenStudy (anonymous):

Using the given zero, find all other zeros of f(x). -2i is a zero of f(x) = x^4 - 32x^2 - 144

OpenStudy (anonymous):

@satellite73 @amistre64

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

since \(-2i\) is a zero, so is its conjugate \(+2i\) and you can now put out two factors \[(x-2i)(x+2i)=x^2+4\]

OpenStudy (anonymous):

that makes \[x^4 - 32x^2 - 144=(x^2+4)(\text{something})\] and you can find the something by division or by thinking

OpenStudy (anonymous):

can Iuse synthetic division? with -4?

OpenStudy (anonymous):

alternatively you can write \[ x^4 - 32x^2 - 144=u^2-32u-144\] where \(u=x^2\) first solve for \(u\) then for \(x\)

OpenStudy (anonymous):

no you cannot

OpenStudy (anonymous):

because it is \(x^2+4\) not \(x+4\) synthetic division will not work here because of the \(x^2\) term

OpenStudy (anonymous):

ohh okay then how would I solve for u and x?

OpenStudy (anonymous):

\[u^2-32u-144=0\\ (u+4)(u-36)=0\\ u=-4,u=36\]

OpenStudy (anonymous):

then \[x^2=-4,x^2=36\] making \[x=\pm2i,x=\pm6\]

OpenStudy (anonymous):

alternatively you can skip the substitution and write \[x^4-32x-144=(x^2+4)(x^2-36)\] and solve that one same thing

OpenStudy (anonymous):

so it's just 2i, 6, -6?

OpenStudy (anonymous):

yes, and also \(-2i\)

OpenStudy (anonymous):

it is a polynomial of degree 4, how many zeros did you expect?